OpenStudy (anonymous):

" @agentx5: No, you both are 6π-π/3. It's true." " @Ishaan94: you can't express pi as rational, i am pretty sure my age is rational " True or False, does 6\(\pi\)-\(\frac{\pi}{3}\) lie on the interval between 17 and 18? ;-)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

my age is fixed fraction. just because 3<pi<4, doesn't make pi someone's age

OpenStudy (anonymous):

Ah but if you're measuring time, or any measurement, do you not cross over that point that represents the irrational at some point?

OpenStudy (anonymous):

What says age can be expressed irrationally?

OpenStudy (anonymous):

In fact... Example: http://www.piday.org/

OpenStudy (anonymous):

it will but like i said i am pretty sure my age is fixed fraction and not pi atm.

OpenStudy (anonymous):

Why do you not celebrate tau day, too?

OpenStudy (anonymous):

i don't even celebrate pi day :( i am not a true mathematician :(

OpenStudy (anonymous):

You are right! Idiotic stuff...

OpenStudy (anonymous):

 I agree your age isn't \(\pi\) yrs., that would make you 3.1415... yrs. old :-D In fact I don't know what you age is, I just am making a mathematics debate in contention that the units don't restrict the numeric value to a integer or repeated decimal. \[\pi \in \mathbb{R}\], agreed?